OpenStudy (anonymous):

bellla is three times as old as edward,jacob is one-half the age of alice. alice is 4 year older than edward bella is 24 years old . who is the youngest

OpenStudy (anonymous):

b =3e j=a/2 a = e+4 b=24 just transcribe the words into eqns. start plugging in, find e from b=3e and b=24. when you have e, find a with a=e+4. continue until you have all of them.

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

a= Alica b = Bella e = Edward j= Jacob whenever you have this type of word problems, show relations with variables and equations...it makes it easier

OpenStudy (anonymous):

when they say someone (let's call them x) is 3times as old as y, the eqn would be : x=3y

OpenStudy (anonymous):

so like how do i find out the awnser

OpenStudy (anonymous):

make mathematical equations (relationships between the variables) from the story they give you. Then, start plugging in your knowns

OpenStudy (anonymous):

do you understand the eqns?

OpenStudy (anonymous):

24=3xE that means E = 8 A = E+4 that means A = 12 J = A/2 that means J = 6 Answer: Bella - 24 Edward - 8 Alice - 12 Jacob - 6

OpenStudy (anonymous):

in other words, jacob is the youngest.

OpenStudy (anonymous):

thank you i get it

OpenStudy (anonymous):

no problem :D